Two objects of mass M connected by a ligft rigid rod are place on a smooth horizontal surface.An impulse J=MV is applied to one of the objects perpendicular to the rod.Calculate the angular accleration of the system just after the impact

Grade:12th pass

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear Student,

Angular momentum impulse given = MV(L/2) = Isystem w
\Rightarrow MV(L/2) = (2ML2/4) w
\RightarrowAngular velocity just after the impact = w = 2V/L

Thanks